# Sparks of Tabular Reasoning via Text2SQL Reinforcement Learning

### v1

@@ Abstract
Empirically, our approach improves performance on standard Text-to-SQL benchmarks and achieves substantial gains on reasoning-intensive datasets such as BIRD and CRT-QA, demonstrating enhanced generalization and interpretability. Specifically, the distilled-quantized LLaMA model achieved a 20% increase in accuracy when trained on Text-to-SQL tasks, while Qwen achieved a 5% increase. These results suggest that SQL can serve not only as a target formalism but also as an effective scaffold for learning robust, transferable reasoning over structured data.
@@ 4.1 Setup
Evaluation Metrics: We employ task-appropriate evaluation metrics for each benchmark. For Text-to-SQL tasks, we report execution accuracy, defined as the exact match between the predicted and reference SQL queries. Due to the partial absence of full database access, we use OpenAI’s o3-mini model as a proxy for execution, assessing query correctness based on structural and semantic alignment. For CRT-QA, where responses may include text spans, numbers, or phrases, o3-mini is again used to evaluate answer correctness.
@@ 4.2 Tabular Aha-Moments
During reinforcement learning with GRPO, we observe instances of emergent tabular reasoning, which we term Tabular Aha-Moments. These moments, inspired by the Aha Moment concept from DeepSeek-R1 [10], occur when the model, provided only a natural language question and schema (but no table content), implicitly reconstructs the structure of the underlying table and uses this to solve the query. An example of this behavior is shown in Figure 4, where the model demonstrates schema-grounded inference without explicit tabular context during the training process.
When evaluated on tabular question answering tasks, the model often invokes SQL-like structures as intermediate reasoning tools—even when SQL output is not required. This is illustrated in Figure 3, where the model constructs an internal SQL representation to derive a binary answer. This reflects a bidirectional inductive bias: the model not only learns to generate SQL from questions but also learns to use SQL representations to support reasoning over tables. These findings highlight the potential for GRPO to induce transferable, structure-aware reasoning in LLMs.
@@
o1 59.1% 20.2% 69.6%
LLaMA Base 37.4% 2% 41.3%
LLaMA SFT 62.1% 1% 33.7%
LLaMA SFT-CoT 56.3% 2% 43.7%
LLaMA SFT-CoT-GRPO 57% 8 % 61.3%
Qwen Base 46.1% 11.9% 49%
Qwen SFT 56.6% 6% 45.3%
Qwen SFT-CoT 53.6% 6.5% 43.2%
Qwen SFT-CoT-GRPO 52.2% 11% 54%
Table 1: Performance comparison of OpenAI o1, the 4-bit quantized version of the distilled Deepseek LLaMA 8B model, and the Qwen-7B-Instruct model (Qwen) evaluated across various datasets. This table compares the performance of untrained models (Base), those supervised fine-tuned on the Clinton Dataset (SFT), models fine-tuned with Chain-of-Thoughts (SFT-CoT) on the Clinton Dataset, and models that have undergone SFT-CoT on the Clinton Dataset and GRPO on the BIRD Dataset. Evaluation scores include execution accuracy determined by an OpenAI o3-mini LLM judge (LLM-EXE), and accuracy determined by an OpenAI o3-mini LLM judge (LLM-ACC).
@@ 4.6 Reward Ablation
Best Reward (LLaMA) 8 57
Best Reward (Qwen) 9 54
Best Two Rewards (LLaMA) 8 56
Best Two Rewards (Qwen) 10 55
Best Four Rewards (LLaMA) 6 64
Best Four Rewards (Qwen) 11 54
Table 2: Ablation study of reward configurations for LLMs trained on the BIRD dataset. The models initially underwent SFT on Chain-of-Thought traces on Clinton, followed by GRPO on BIRD, where specific reward functions were applied. Performance is evaluated across the best reward configurations (best one, two, and four rewards) for each model. Evaluation scores include execution accuracy (LLM-EXE) and accuracy (LLM-ACC) assessed by an LLM judge. The table compares the models’ performance across different configurations of rewards and tasks.
@@ 5 Conclusion
In conclusion, our experiments demonstrate that integrating reinforcement learning with a supervised pretraining phase significantly enhances the model’s ability to reason over tabular data. Notably, the distilled quantized LLaMA model achieved a 20% performance improvement on the CRT-QA dataset, while the Qwen model saw a 5% increase, underscoring the efficacy of our two-stage framework in optimizing SQL execution and fostering transferable reasoning for complex question answering tasks. These findings suggest that SQL serves not only as a task-specific format but also as a foundational scaffold for developing robust tabular reasoning skills in LLMs.

### v2

@@ Abstract
Empirically, our approach improves performance on standard Text-to-SQL benchmarks and achieves substantial gains on reasoning-intensive datasets such as BIRD (text-to-SQL) and CRT-QA (tabular QA), demonstrating enhanced generalization and interpretability. Specifically, the distilled-quantized LLaMA model achieved a relative 33.9% increase in tabular reasoning accuracy when trained on Text-to-SQL tasks, while Qwen achieved a relative 14.5% increase. These results suggest that SQL can serve not only as a target formalism but also as an effective scaffold for learning robust, transferable reasoning over structured data.
@@ 4.1 Setup
Evaluation Metrics: We employ task-appropriate evaluation metrics for each benchmark. For Text-to-SQL tasks, we report execution accuracy, defined as the exact match between the predicted and reference SQL queries. Due to the partial absence of full database access, we use OpenAI’s o3-mini model as a proxy for execution, assessing query correctness based on structural and semantic alignment. While o3-mini offers a scalable and semantically sensitive evaluation mechanism, it is not a substitute for full SQL execution against a live database. Future work will incorporate actual DBMS-based evaluation to better ground model performance in execution fidelity.
@@ 4.2 Tabular Aha-Moments
During reinforcement learning with GRPO, we observed qualitatively interesting behaviors that resemble the Aha Moment described in DeepSeek-R1 [10]. Specifically, the model occasionally demonstrated schema-aware reasoning: it was able to infer structural relationships and generate meaningful outputs based solely on a natural language question and the corresponding database schema—without access to the underlying table content.
An example of this behavior is shown in Figure4, where the model employs its understanding of tables to produce accurate SQL queries. Additionally, when evaluated on tabular question answering tasks, the model frequently utilizes SQL-like structures as intermediate reasoning tools, even when SQL output is not necessary. This is illustrated in Figure 3, where the model constructs an internal SQL representation to derive a binary answer.
These instances suggest that GRPO may contribute to the development of structure-aware reasoning capabilities in large language models. However, we treat these findings as exploratory and leave rigorous measurement of such schema-grounded generalization to future work.
@@
o1 59.1% 25.4% 69.6%
LLaMA Base 37.4% 2.8% 41.3%
LLaMA SFT 62.1% ↑ +24.7 1.3% 33.7%
LLaMA SFT-CoT 56.3% 4.7% 43.7%
LLaMA SFT-CoT-GRPO 57.0% 8% ↑ +5.2 55.3% ↑ +14.0
Qwen Base 46.1% 11.9% 49.0%
Qwen SFT 56.6% ↑ +10.5 6.1% 45.3%
Qwen SFT-CoT 53.6% 9.5% 45.2%
Qwen SFT-CoT-GRPO 52.2% 15.5% ↑ +3.6 56.1% ↑ +7.1
Table 1: Performance on Tabular QA and Text-to-SQL.Performance comparison of OpenAI o1, the 4-bit quantized version of the distilled Deepseek LLaMA 8B model, and the Qwen-7B-Instruct model (Qwen) evaluated across various datasets. This table compares the performance of untrained models (Base), those supervised fine-tuned on the Clinton Dataset (SFT), models fine- tuned with Chain-of-Thoughts (SFT-CoT) on the Clinton Dataset, and models that have undergone SFT-CoT on the Clinton Dataset and GRPO on the BIRD Dataset. Notable improvements are seen with LLaMA SFT (+24.7%) and Qwen SFT (+10.5%) on the Clinton dataset, demonstrating the effectiveness of supervised fine-tuning. Our main result is that SFT-CoT-GRPO methodologies enhance performance on CRT-QA (tabular reasoning) with gains of +14.0% and +7.1%.
@@ 4.6 Reward Ablation
Best Reward (LLaMA) 8.2% 51.7%
Best Reward (Qwen) 9.1% 54.3%
Best Two Rewards (LLaMA) 8.6% 54.8%
Best Two Rewards (Qwen) 10.8% 55.0%
Best Four Rewards (LLaMA) 8.0% 55.3%
Best Four Rewards (Qwen) 15.5% 56.1%
Table 2: Ablation study of reward configurations. The models initially underwent SFT on Chain-of-Thought traces on Clinton, followed by GRPO on BIRD, where specific reward functions were applied. Performance is evaluated across the best reward configurations: the best reward is derived from string matching, the best two consist of both string matching and evaluations by an LLM judge, and the best four are based on a combination of string matching, component-level matching, execution-based reward, and evaluations by an LLM judge. Evaluation scores include execution accuracy (LLM-EXE) and accuracy (LLM-ACC) assessed by an LLM judge.
@@ 5 Conclusion
Our experiments demonstrate that treating Text-to-SQL as a structured reasoning task—rather than just query generation—significantly enhances LLMs’ ability to understand tabular data. We use a two-stage framework:
  1. 1.
    Synthetic Chain-of-Thought Supervision: generates detailed, clause-level reasoning traces to teach models how to traverse, filter, and aggregate table fields.
  2. 2.
    Reinforcement Learning with GRPO: leverages multiple reward signals to align SQL execution accuracy with broader, transferable reasoning behaviors.
Our work goes beyond execution optimization by integrating multi-dimensional reward signals and leveraging the effects of GRPO for broader tabular reasoning tasks, such as question answering (QA). In this way, GRPO is not just an optimizer but a vehicle for generalization and reasoning alignment. The distilled and quantized LLaMA model achieves a relative 33.9% increase, while the Qwen model sees a relative 14.5% improvement on the tabular reasoning CRT-QA benchmark.
These results confirm that SQL serves not only as a task-specific formalism but also as an effective scaffold for teaching robust, transferable reasoning over structured data. Moving forward, we plan to scale this method to larger models, explore richer reward configurations, and extend the framework to other formal languages and real-world datasets, aiming to build more interpretable and capable systems for structured data interaction.